OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 5x - 3y = 29 and x - 4y = 16

OpenStudy (compassionate):

Hello, Suzumiya. 5x - 3y = 29 x - 4y = 16 Do you know the first step?"

OpenStudy (anonymous):

no

OpenStudy (compassionate):

First step is getting the x or y terms to cancel. Lets get the y terms to cancel 4(5x - 3y = 29) 3(x - 4y = 16) Now I would like for you to multiple each term by 4 and 3, just like in the equation above.

OpenStudy (anonymous):

20x-12y=116 3x-12y=48

OpenStudy (compassionate):

You just keep posting questions. I would like you to try and solve it on your own, then I'll go back and correct you mistakes.

OpenStudy (anonymous):

(4, 3) (-4, 3) (4, -3) no solution infinitely many solutions my answer i don't know what to do :(

OpenStudy (compassionate):

My, my. Didn't you learn this in school? http://prntscr.com/13yos5 Use a graphing calculator?

OpenStudy (anonymous):

i learn better by some one show me how to do the steps